# SubGrapher: Visual Fingerprinting of Chemical Structures

### v1

@@ 4 Abstract
Automatic extraction of chemical structures from scientific literature plays a crucial role in accelerating research across fields ranging from drug discovery to materials science. Patent documents, in particular, contain molecular information in visual form, which is often inaccessible through traditional text-based searches. In this work, we introduce SubGrapher, a method for the visual fingerprinting of chemical structure images. Unlike conventional Optical Chemical Structure Recognition (OCSR) models that attempt to reconstruct full molecular graphs, SubGrapher focuses on extracting molecular fingerprints directly from chemical structure images. Using learning-based instance segmentation, SubGrapher identifies functional groups and carbon backbones, constructing a substructure-based fingerprint that enables chemical structure retrieval. Our approach is evaluated against state-of-the-art OCSR and fingerprinting methods, demonstrating superior retrieval performance and robustness across diverse molecular depictions. The dataset, models, and code will be made publicly available.
@@ 5 Introduction
Knowledge in chemistry is spread across structured databases and unstructured sources such as scientific journals, patent documents, books, and corporate documents. Integrating molecular structures, properties, and synthesis data into a unified collection would significantly accelerate research in chemistry and materials science [1]. Converting unstructured documents to machine readable format enables searching within document collections and training foundational models at larger scales [2, 3]. A key challenge in this process is extracting molecular structures from patent documents, where they are primarily represented as images rather than machine-readable text. Molecular databases such as PatCID [4] and SureChEMBL [5] aim to address this challenge by employing document ingestion pipelines that integrate page segmentation, chemical image classification, and Optical Chemical Structure Recognition (OCSR).
@@
Despite advances, OCSR methods face challenges with variations in chemical drawing conventions or degraded image quality. Additionally, certain chemical illustrations cannot be represented as SMILES, particularly those containing non-standard graphical elements (see row 3 in Figure 4) or Markush structures [14]. This limitation is especially significant in patent analysis, where Markush structures are widely used to define broad molecular classes. Moreover, full molecular reconstruction is not always necessary for applications such as database searching or molecular property prediction. In many cases, users are more interested in identifying molecules with specific substructures rather than their complete structures. Furthermore, predictive models often rely on molecular fingerprints derived from SMILES [15], making SMILES an intermediate representation rather than a final output.
Figure 1: SubGrapher extracts a molecular fingerprint from a chemical structure image in a document. Our approach identifies functional groups and carbon backbones in chemical structure images. These substructures are then combined based on their connectivity to create a substructure-graph. Finally, this graph is converted to a fingerprint enabling substructure search, similarity search or any downstream predictive task.
@@
Molecular fingerprints are vectorized representations of molecular structures, commonly used for similarity searches in databases and predictive tasks such as property prediction or drug target identification [16]. Substructure-based fingerprints, such as MACCS [17], PubChem Fingerprints [18], RDKit Daylight [19], and Extended-Connectivity Fingerprints (ECFP) [20], encode the presence or absence of predefined molecular fragments. Among these substructures, functional groups [21] form a key subset due to their well-defined chemical properties. Besides, molecular fingerprints can also be generated using probabilistic methods like MHFP [22] or learned from large datasets using models such as MoLFormer [23].
@@
Integrating OCSR and fingerprinting into a single-step process by directly recognizing functional groups from images would efficiently facilitate the extraction of molecular information from literature. Recently, Fan, et al., [24] explored the recognition of a limited subset of functional groups from images as an auxiliary task to enhance OCSR performance. Additionally, other studies [25, 26] have investigated the use of molecule images to learn molecular features used for predictive tasks. However, none of these studies have combined visual recognition of functional groups with visual fingerprinting into a unified approach.
@@
In this work, we introduce SubGrapher for the visual fingerprinting of chemical structure images, illustrated in Figure 1. First, SubGrapher uses segmentation models to identify functional groups and carbon backbones in images. Second, SubGrapher creates a substructure-graph based on the connectivity of these substructures. Finally, this graph is converted to a molecular fingerprint. On the one hand, our fingerprint allows to search for molecules containing functional groups and carbon backbones of interest. By defining key chemical properties, functional groups are natural targets for molecular searches, while carbon backbones link these functional groups. On the other hand, SubGrapher’s fingerprint can be used for approximate search of exact chemical structures.
@@ 6 Method
We introduce SubGrapher, a model designed for the visual fingerprinting of chemical structure images. Its architecture is illustrated in Figure 2. SubGrapher employs substructure segmentation models to detect and identify molecular substructures, specifically functional groups and carbon backbones. These detected substructures are then assembled into a substructure-graph, which is subsequently converted into a molecular fingerprint.
Figure 2: SubGrapher architecture. The SubGrapher model detects functional groups and carbon backbones using instance segmentation networks. The identified substructures are combined into a substructure-graph, which is then converted to a matrix fingerprint.
@@ 6.1 Substructure Segmentation Model
To detect molecular substructures from images, SubGrapher employs two segmentation networks, as illustrated in Figure 2. The first network detects 1,534 expert-defined functional groups, ensuring broad coverage of molecular diversity. The second network identifies 27 distinct carbon backbone patterns, complementing the functional group detection by capturing molecular regions not already assigned to any functional groups. By ensuring that substructures overlap appropriately and span the entire molecule, SubGrapher constructs a more characteristic and informative fingerprint in later stages.
Substructures are detected using mask-based segmentation rather than relying solely on bounding-box detection. This approach provides fine-grained supervision during training, leading to improved accuracy and robustness in molecular substructure recognition. Additionally, our framework enables training substructure recognition directly from page images, eliminating the need for segmentation in document-to-chemical-structure pipelines [4]. In the implementation, the segmentation models are Mask-RCNN [27] models.
@@ 6.2 Substructure-Graph Construction
The detected substructures are used to construct a substructure-molecular graph, where nodes represent the identified substructures and edges correspond to their intersections.
Substructures are considered intersected if their predicted bounding boxes overlap. To enhance connectivity, each detected box is expanded by a fixed amount. This expansion ensures that adjacent substructures sharing a single atom are linked. Additionally, substructures that are not directly connected in the molecular graph may still form connections in the substructure-graph if they are spatially close in the depicted molecular conformation. This added connectivity incorporates positional information, improving the representativity of the substructure graph. In the implementation, the expansion of each bounding box is set to 10% of the diagonal length of the smallest detected box in the image. This constraint prevents excessive expansion, ensuring a reasonable balance that avoids artificially distant intersections.
@@ 6.3 Substructure-based Visual Molecular Fingerprint
To describe molecular structures, SubGrapher constructs a visual molecular fingerprint by combining detected substructures. This fingerprint encodes the occurrences of functional groups and carbon segments, as well as their distances.
We introduce the substructure-based visual molecular fingerprint (SVMF). For a given molecule $m$, SVMF$(m)$ is defined as an upper triangular matrix:
$\text{SVMF}(m)=\begin{bmatrix}&\vspace{-2mm}\ddots&&&&&&&\\ &&&\hskip 5.69054ptf_{i,i}&\hskip 5.69054pt\dots&\hskip 5.69054ptf_{i,j}&&&\\ &&&&\vspace{-2mm}\ddots&\hskip 5.69054pt\vdots&&&\\ &&&{\text{0}}&&\hskip 5.69054ptf_{j,j}&&&\\ &&&&&&\ddots&&\\ \end{bmatrix}$ (1)
[…]
$f_{i,i}=h_{1}.n_{i}$ (2)
where $n_{i}$ is the number of instances of substructure $i$ in the molecule, and $h_{1}$ is a model hyperparameter. The off-diagonal elements, $f_{i,j}$, represent the intersection coefficient between substructures $i$ and $j$:
$f_{i,j}=\sum_{\alpha=1}^{n_{i}}\sum_{\beta=1}^{n_{j}}h_{2}\circ d(i_{\alpha},j_{\beta})$ (3)
where $\circ$ is the composition operator, $d(i_{\alpha},j_{\beta})$ denotes the distance between instances $i_{\alpha}$ and $j_{\beta}$, and $n_{i}$ and $n_{j}$ are the number of instances of the substructures $i$ and $j$. The distance $d(i_{\alpha},j_{\beta})$ is defined as the smallest number of substructures required to construct a path between instances $i_{\alpha}$ and $j_{\beta}$. The function $h_{2}$ is an hyper-parameter of the model controlling the number of non-zero intersection coefficients, and ultimately the fingerprint compression.
In the implementation, $h_{1}$ is set to 10, while $h_{2}(d)$ takes values of $2$, $2/4$, $2/16$, and $2/256$ for distances $d=1,2,3,4$, respectively. If $d>4$, the intersection coefficient is set to zero. For carbon chain instances, intersection coefficients are divided by $2$ to assign greater significance to functional groups. This prioritization reflects the intuition that functional groups are the most important parts of the molecule. The SVMF matrix can be transformed into a one-dimensional fingerprint by first converting all values to integers using a fixed precision cutoff and then separately encoding active matrix positions and their corresponding values.
@@ 6.4 Molecule Retrieval
where $\parallel$.$\parallel$ . ∥ denotes the Euclidian norm. This similarity measure enables molecular searches in large datasets. Given a query molecule, users can retrieve similar structures. Further details are in the Results section below.
6.5 Substructures Coverage
Figure 3: Substructure examples. Examples of functional groups recognized by SubGrapher.
Here, we describe the strategy used to select the 1,534 functional groups and 27 carbon backbone patterns.
Our functional groups are defined as substructures containing at least one heteroatom and having for attachment points single bonds connected to carbon atoms. These substructures are manually designed starting from chemically logical chains of up to five atoms selected from C, O, S, N, B, or P. Each heteroatom in these initial chains can be expanded with chemically relevant subgroups containing atoms C, H, O, S, or N. For instance, if sulfur is present in the primary chain, we also consider replacing it with SO and SO2. To refine the selection, we perform substructure searches in PubChem for candidate ‘functional group family’ (i.e., chemically similar groups). Families with fewer than approximately 1,000 occurrences in PubChem were not further expanded. Some example functional groups are illustrated in Figure 3.
For carbon backbones, we include standard combinations of three to six carbon atoms with single, double, or triple bonds, as well as common three- to six-membered rings. If a carbon backbone substructure is fully contained within another substructure, we retain only the larger one to minimize redundancy. We defined functional groups and carbon backbones substructures to ensure that they are sufficiently overlapping. This allows our molecular fingerprint to remain distinctive and applicable to a wide range of organic compound.
@@ 6.6 Training Dataset Generation
Datasets of molecular images extracted from real documents are scarce. Especially, there is no dataset which includes mask annotations. To address this limitation, we use the synthetic data generation pipeline proposed in MolDepictor [10], which generates a diverse set of molecular depictions. This pipeline selects molecular SMILES from PubChem and renders them using the RDKit library [19].
@@ 7 Results and Discussion
In this section, we perform experiments to evaluate SubGrapher for substructure detection and molecular fingerprinting. Our method is evaluated against state-of-the-art OCSR models and SMILES-based fingerprinting methods. Moreover, we analyze the strengths and weaknesses of SubGrapher on qualitative examples.
@@ 7.1.1 Datasets and Metrics
To compare our method with state-of-the-art approaches, we evaluate it on three benchmarks: JPO [30], a subset of USPTO-10K-L [10], and USPTO-Markush [31]. These datasets contain pairs of chemical images and molecular structures stored as MOL files [32].
JPO consists of 341 molecule images extracted from patent documents published by the Japanese Patent Office [33]. This dataset is challenging due to non-standard drawing conventions and low image quality. Molecules containing abbreviations were removed from the standard JPO benchmark set. USPTO-10K-L contains 10,000 images of large molecules with over 70 atoms and no abbreviations. We evaluate our method using the first 1,000 images of this benchmark. USPTO-Markush includes 74 images of Markush structures from patent documents published by the United States Patent and Trademark Office (USPTO [34]).
@@
To evaluate performances, we compute the Substructure F1-score (S-F1) and the Molecule Exact Match (M-EM). The substructure F1-score measures recall (the proportion of ground-truth substructures correctly identified by predictions) and precision (the proportion of predicted substructures that correctly match the ground truth) for substructure detection. The molecule exact match is the percentage of molecules where all substructures are perfectly detected, i.e., the S-F1 equals one. Since OCSR models such as OSRA and DECIMER do not predict atoms positions, our evaluation metrics consider only the presence or absence of substructures, not their exact location.

### v2

@@ 4 Abstract
Automatic extraction of molecules from scientific literature plays a crucial role in accelerating research across fields ranging from drug discovery to materials science. Patent documents, in particular, contain molecular information in visual form, which is often inaccessible through traditional text-based searches. In this work, we introduce SubGrapher, a method for the visual fingerprinting of molecule and Markush structure images. Unlike conventional Optical Chemical Structure Recognition (OCSR) models that attempt to reconstruct full molecular graphs, SubGrapher focuses on extracting fingerprints directly from images. Using learning-based instance segmentation, SubGrapher identifies functional groups and carbon backbones, constructing a substructure-based fingerprint that enables the retrieval of molecules and Markush structures. Our approach is evaluated against state-of-the-art OCSR and fingerprinting methods, demonstrating superior retrieval performance and robustness across diverse molecule and Markush structure depictions. The benchmark datasets, models, and inference code are publicly available.
4.1 Scientific contribution statement
SubGrapher introduces a novel approach to convert molecule and Markush structure images directly into fingerprints in a single step, bypassing traditional SMILES or graph reconstruction. It outperforms existing OCSR and fingerprinting methods for substructure detection and structure retrieval across diverse datasets, including Markush structure images.
@@ 5 Introduction
Knowledge in chemistry is spread across structured databases and unstructured sources such as scientific journals, patent documents, books, and corporate documents. Integrating molecular structures, properties, and synthesis data into a unified collection would significantly accelerate research in chemistry and materials science [1]. Converting unstructured documents to machine-readable formats enables searching within document collections and training foundational models at larger scales [2, 3, 4]. A key challenge in this process is extracting molecular structures from patent documents, where they are primarily represented as images rather than machine-readable text. Molecular databases such as PatCID [5] and SureChEMBL [6] aim to address this challenge by employing document ingestion pipelines that integrate page segmentation, chemical image classification, and Optical Chemical Structure Recognition (OCSR).
@@
Despite advances, OCSR methods face challenges with variations in drawing conventions or degraded image quality. Additionally, certain chemical illustrations cannot be represented as SMILES, especially depictions with non-standard graphical elements (see row 3 in Figure 4) or Markush structures [15]. This limitation is especially significant in patent analysis, where Markush structures are widely used to define broad molecular classes. Moreover, full molecular reconstruction is not always necessary for applications such as database searching or molecular property prediction. In many cases, users are more interested in identifying molecules with specific substructures rather than their complete structures. Furthermore, predictive models often rely on molecular fingerprints derived from SMILES [16], making SMILES an intermediate representation rather than a final output.
@@
Molecular fingerprints are vectorized representations of molecular structures, commonly used for similarity searches in databases and predictive tasks such as property prediction or drug target identification [17]. Structural key fingerprints, such as MACCS [18] and PubChem fingerprints [19], encode as a binary vector the presence or absence of predefined molecular fragments. Among these fragments, functional groups [20] form a key subset due to their well-defined chemical properties. To improve generalization, other approaches avoid relying on predefined fragment libraries. Instead, they generate fragments directly from the molecular graph, which are then hashed and folded into bit vectors. For example, Daylight-style fingerprints [21] generate fragments by enumerating linear atom–bond paths, while Extended-Connectivity Fingerprints (ECFP) [22] generate fragments by enumerating circular neighborhoods around atoms. In addition, molecular fingerprints can also be generated using probabilistic techniques like MinHash fingerprints (MHFP) [23], or learned from large datasets using deep learning models such as MoLFormer [24].
@@
Integrating OCSR and fingerprinting into a single-step process by directly recognizing functional groups from images would efficiently facilitate the extraction of molecular information from literature. Recently, Fan, et al., [25] explored the recognition of a limited subset of functional groups from images as an auxiliary task to enhance OCSR performance. Additionally, other studies [26, 27] have investigated the use of molecule images as a way to learn molecular features for downstream predictive tasks. However, none of these studies have combined visual recognition of functional groups with visual fingerprinting into a unified approach.
@@
In this work, we introduce SubGrapher for the visual fingerprinting of molecule and Markush structure images, illustrated in Figure 1. First, SubGrapher uses segmentation models to identify functional groups and carbon backbones in images. Second, SubGrapher creates a substructure-graph based on the connectivity of these substructures. Finally, this graph is converted to a count-based continuous fingerprint. On the one hand, our fingerprint allows to search for molecules containing functional groups and carbon backbones of interest. By defining key chemical properties, functional groups are natural targets for molecular searches, while carbon backbones link these functional groups. On the other hand, SubGrapher’s fingerprint can be used for molecule and Markush structure retrieval.
Figure 1: SubGrapher extracts a fingerprint from a molecular or Markush structure image in a document. Our approach identifies functional groups and carbon backbones in images. These substructures are then combined based on their connectivity to create a substructure-graph. Finally, this graph is converted to a fingerprint enabling substructure search, molecule and Markush structure retrieval, or any downstream predictive task.
@@ 6 Method
We introduce SubGrapher, a model designed for the visual fingerprinting of molecule and Markush structure images. Its architecture is illustrated in Figure 2. SubGrapher employs substructure segmentation models to detect and identify molecular substructures, specifically functional groups in the organic chemistry domain and carbon backbones. These detected substructures are then assembled into a substructure-graph and converted into a count-based continuous fingerprint.
Figure 2: SubGrapher architecture. The SubGrapher model detects functional groups and carbon backbones using instance segmentation networks. The identified substructures are combined into a substructure-graph, which is then converted to a matrix fingerprint. Finally, the matrix is stored as a compressed vector. The resulting fingerprint is count-based, as each coefficient depends on the number of substructure occurrences, and continuous, since the values are real numbers.
@@ 6.1 Substructure Segmentation Model
To detect molecular substructures from images, SubGrapher employs two segmentation networks, as illustrated in Figure 2. The first network detects 1534 expert-defined functional groups, defined in the Substructures Coverage section below. The second network identifies 27 distinct carbon backbone patterns, complementing the functional group detection by capturing molecular regions not already assigned to any functional groups. By ensuring that substructures overlap appropriately and span the entire molecule, SubGrapher constructs a more characteristic and informative fingerprint in later stages.
Substructures are detected using mask-based segmentation rather than relying solely on bounding-box detection. This approach provides fine-grained supervision during training, leading to improved accuracy and robustness in molecular substructure recognition. In the implementation, the segmentation models are Mask-RCNN [28] models.
6.2 Substructures Coverage
Figure 3: Substructure examples. Examples of functional groups and carbon backbones recognized by SubGrapher.
Here, we describe the strategy used to select the 1534 functional groups and 27 carbon backbone patterns. Each substructure is manually defined and annotated with a SMILES, SMARTS [29], and a descriptive name.
Our functional groups are defined as substructures containing at least one heteroatom and having for attachment points single bonds connected to carbon atoms. These substructures are manually designed starting from chemically logical chains of up to five atoms selected from C, O, S, N, B, or P. Each heteroatom in these initial chains can be expanded with chemically relevant subgroups containing atoms C, H, O, S, or N. For instance, if sulfur is present in the primary chain, we consider replacing it with SO and SO2. To refine the selection, we perform substructure searches in PubChem [19] for candidate ‘functional group family’ (i.e., chemically similar groups). Families with fewer than approximately 1000 occurrences in PubChem were not included. In addition, we added a set of halogen substituents and organometallic groups relevant to extreme ultraviolet (EUV) photoresists to the functional group list. Examples of functional groups are shown in Figure 3, and a visualization of their diversity is illustrated in Supplementary Figure 1. Our list is among the most extensive available in the open-source domain [20, 30, 31, 32], yet its scope remains limited to widely occurring substructures and to organic chemistry.
For carbon backbones, we include standard combinations of three to six carbon atoms with single, double, or triple bonds, as well as common three- to six-membered rings. If a carbon backbone substructure is fully contained within another substructure, we retain only the larger one to minimize redundancy. We defined functional groups and carbon backbones substructures to ensure that they are sufficiently overlapping. As discussed in the following section, this overlap will allow to create a more distinctive fingerprint. Additionally, in Supplementary Note 1, we support the relevance of our substructures by evaluating their coverage on 122M molecules from PubChem.
@@ 6.3 Substructure-Graph Construction
In this section, we present how the detected substructures are used to construct a substructure-graph. The nodes of the substructure-graph represent the identified substructures and its edges correspond to the intersections between these substructures.
Substructures are considered intersected if their predicted bounding boxes overlap. To enhance connectivity, each detected box is expanded by a margin. This expansion ensures that adjacent substructures sharing a single atom are linked. Additionally, substructures that are not directly connected in the molecular graph may still form connections in the substructure-graph if they are spatially close in the depicted molecular conformation. This added connectivity incorporates positional information, improving the representativity of the substructure graph. In the implementation, the expansion of each bounding box is set to 10% of the diagonal length of the smallest detected box in the image. This expansion influences which groups are overlapping. Supplementary Table 1 shows that the similarity between resulting fingerprints is minimally sensitive to this parameter. A relatively small expansion is preferred to limit the overall size of the fingerprint.
@@ 6.4 Substructure-based Visual Molecular Fingerprint
To describe molecules and Markush structures, SubGrapher constructs a visual molecular fingerprint by combining detected substructures. This fingerprint encodes the occurrences of functional groups and carbon segments, as well as their distances.
We introduce the Substructure-based Visual Molecular Fingerprint (SVMF). For a given molecule $m$, SVMF(m)(m) is defined as an upper triangular matrix:
$\text{SVMF}(m)=\begin{bmatrix}&\vskip-5.69054pt\ddots&&&&&&&\\ &&&\hskip 5.69054ptf_{i,i}&\hskip 5.69054pt\dots&\hskip 5.69054ptg_{i,j}&&&\\ &&&&\vskip-5.69054pt\ddots&\hskip 5.69054pt\vdots&&&\\ &&&{\text{0}}&&\hskip 5.69054ptf_{j,j}&&&\\ &&&&&&\ddots&&\\ \end{bmatrix}$ (1)
$\text{SVMF(m)}\in\mathbb{R}^{n\times n}$, with $n$ is the total number of substructures considered. Here, $n$ is equal to 1561, including both functional groups and carbon backbone substructures. The diagonal elements $f_{i,i}$ represent the substructure coefficients and are defined as:
$f_{i,i}=h_{1}.n_{i}+g_{i,i}$ (2)
where $n_{i}$ is the number of instances of substructure $i$ in the molecule, $h_{1}$ is a model hyperparameter and $g_{i,i}$ is defined as below. The off-diagonal elements $g_{i,j}$ represent the intersection coefficient between substructures $i$ and $j$:
$g_{i,j}=\sum_{\alpha=1}^{n_{i}}\sum_{\beta=1}^{n_{j}}h_{2}\circ d(i_{\alpha},j_{\beta})$ (3)
where $\circ$ is the composition operator, $d(i_{\alpha},j_{\beta})$ denotes the distance between instances $i_{\alpha}$ and $j_{\beta}$, and $n_{i}$ and $n_{j}$ are the number of instances of the substructures $i$ and $j$. The distance $d(i_{\alpha},j_{\beta})$ is defined as the smallest number of substructures required to construct a path between instances $i_{\alpha}$ and $j_{\beta}$. The function $h_{2}$ is an hyper-parameter of the model controlling the number of non-zero intersection coefficients, and ultimately the fingerprint compression. The SVMF is count-based, as each coefficient depends on the number of substructure occurrences, and continuous, since the values are real numbers.
In the implementation, $h_{1}$ is set to 10, while $h_{2}(d)$ takes values of $2$, 22, $2/4$, $2/16$, and $2/256$ for distances $d=0,1,2,3,4$, respectively. If $d>4$, the intersection coefficient is set to zero. For carbon chain instances, intersection coefficients are divided by $2$ to assign greater significance to functional groups. This prioritization reflects the intuition that functional groups are the most important parts of the molecule. The SVMF, which has a dimension $1561\times 1561=1\times 2436721$, is not stored in its dense form. Instead, only the non-zero coefficients and their corresponding values are encoded in a compressed vector. The average number of non-zero coefficients for the evaluated fingerprints is reported in the Supplementary Table 2. For the SVMF, the average proportion of non-zero coefficients is 0.0010.001%.
@@ 6.5 Molecule Retrieval
where $\parallel.\parallel$ denotes the Euclidean norm. This measure enables searching in large datasets by comparing a query molecule or Markush structure with others to retrieve structurally similar structures. Further details are in the Results section below.
@@ 6.6 Training Dataset Generation
Datasets of molecule and Markush structure images extracted from real documents are scarce. Especially, there is no dataset which includes mask annotations. To address this limitation, we use the synthetic data generation pipeline introduced in MolDepictor [11], which generates a diverse set of molecular depictions. This pipeline selects molecular SMILES from PubChem and renders them using the RDKit library [33].
@@ 7 Results and Discussion
In this section, we perform experiments to evaluate SubGrapher for substructure detection and fingerprinting. Our method is evaluated against state-of-the-art OCSR models and SMILES-based fingerprinting methods.
@@ 7.1.1 Datasets and Metrics
To compare our method with state-of-the-art approaches for substructure detection, we evaluate it on three benchmarks: JPO [36], a subset of USPTO-10K-L [11], and USPTO-Markush [37]. These datasets contain pairs of images and structures stored as MOL files [38].
JPO consists of 341 molecule images extracted from patent documents published by the Japanese Patent Office [39]. This dataset is challenging due to non-standard drawing conventions and low image quality. Molecules containing abbreviations were removed from the standard JPO benchmark set as SubGrapher is currently unable to recognize functional groups in abbreviations. USPTO-10K-L contains 10000 images of large molecules with over 70 atoms and no abbreviations. We evaluate our method using the first 1000 images of this benchmark. USPTO-Markush includes 74 images of Markush structures from patent documents published by the United States Patent and Trademark Office (USPTO [40]).
@@
To evaluate performances, we compute the Substructure F1-score (S-F1) and the Molecule Exact Match (M-EM). The substructure F1-score measures recall (the proportion of ground-truth substructures correctly identified by predictions) and precision (the proportion of predicted substructures that correctly match the ground truth) for substructure detection. The molecule exact match is the percentage of molecules where all substructures are perfectly detected, i.e., the S-F1 equals one.
SubGrapher is compared with three OCSR methods used for substructure detection: MolGrapher, OSRA, and DECIMER. These methods have the advantage of being more general, as converting images to SMILES enables the detection of any functional group, including those outside the scope of our evaluation set. However, OSRA and DECIMER do not predict atoms positions, therefore our evaluation metrics consider only the presence or absence of substructures, not their exact location.